OpenStudy (anonymous):

multiply or divide as indicated and reduce to lowest terms where appropriate: 2x-4/5x X 15x/x-2 multiply or divide as indicated and reduce to lowest terms where appropriate: 2x-4/5x X 15x/x-2 @Mathematics

OpenStudy (anonymous):

how the hell?!

OpenStudy (anonymous):

I assume that is \[\frac{2x-4}{5x}\frac{15x}{x-2} \] right?

OpenStudy (anonymous):

You can rewrite that like this... \[\frac{2(x-2)}{5x}\frac{3(5x)}{x-2} \] which would be six, no?

OpenStudy (anonymous):

correct how did u write it like that? i cant write it like that in the question box

OpenStudy (anonymous):

but i dont know how you are even re writing it

OpenStudy (anonymous):

It's called LaTeX, don't worry about it. And just look at it, I dunno what else to tell you really. 2x-4 = 2(x-2), right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So, the (x-2)'s cancel, and the 5x's cancel, leaving you with 2*3 = 6.

OpenStudy (anonymous):

ohhhhh okay thank you!

OpenStudy (anonymous):

so what about -8 X n/-3msquared

OpenStudy (anonymous):

you can cancel the negative signs.

OpenStudy (anonymous):

this is such hard math

OpenStudy (anonymous):

i dont know how to calculate algebraic functions

OpenStudy (anonymous):

Such as?

OpenStudy (anonymous):

i dont know to write it

OpenStudy (anonymous):

(5/pqSQUARED) ( -3pCUBEDr/10q)

OpenStudy (anonymous):

\[\frac{5}{pq^2}\frac{-3p^3r}{10q}\] ?

OpenStudy (anonymous):

how do u do that?

OpenStudy (anonymous):

You'll notice the "Equation" button just below the text box.

OpenStudy (anonymous):

i see it but i havent figured out how to work it

OpenStudy (anonymous):

It's a document typesetting language called LaTeX. If you can't work the buttons don't worry about it because what I just typed looks like this: "\frac{5}{pq^2} \frac{-3p^3r}{10q} "

OpenStudy (anonymous):

I understood what you meant well enough when you just typed it.

OpenStudy (anonymous):

okie doke

OpenStudy (anonymous):

And to answer your question, that would be \[\frac{-15p^3r}{10pq^3} = \frac{-3p^2r}{2q^3}\]

OpenStudy (anonymous):

can you explain how you got that?

OpenStudy (anonymous):

I can tell you that when you multiply fractions, you multiply the numerators and denominators together, and that 15 and 10 share a common factor of 5 so that reduces to 3/2 when you cancel it, and p^3 / p = p^2, but besides that, I apologize, I do not know how to explain it further.

OpenStudy (anonymous):

thanks i just not that good at math